A clerk can complete a job in 4 hours. Another clerk can do the same job in 2 hours. How long will it take them working together?
This is what I got and wanted to see if it was right? 1/4x + 1/2x=1
then i put 1/4 + 2/4 =4t
3x/3 = 4/3 = 1 1/3 hours

Answer by checkley71(8403)

A MORE DIRECT APPROACH TO THESE TYPE OF BROBLEMS IS THE PRODUCT DIVIDED BY THE SUM OF THE 2 TIMES.
4*2/(4+2)=8/6=4/3=1 1/3 ANSWER.